OpenStudy (anonymous):

simplify 2+17x-5x+9 a. 19x + a b. 12x + 11 c. 22x + 11 d. 15x - 4

OpenStudy (anonymous):

fan and medal

OpenStudy (anonymous):

@ballerinaxbb

OpenStudy (ghostgate):

Well add and subtract like terms, for instance: 17x - 5x = ?x 9 + 2 = ? I'm pretty sure you can reduce it from there. I hope this helps in someway! Have a great day! ;) {---Ghostgate---}

OpenStudy (anonymous):

no not much @Ghostgate

OpenStudy (ghostgate):

Alright, what is 9 + 2?

OpenStudy (ghostgate):

And then what is 17 - 5?

OpenStudy (anonymous):

so it's b?

OpenStudy (ghostgate):

Yes, because 17x - 5x = 12x and 9 + 2 = 11, leaving you with: 12x + 11

OpenStudy (anonymous):

thanks

OpenStudy (ghostgate):

No problem, I'm glad I could help!